Write the missing number for:
? divided by 35 = 3 remainder24

Answer by checkley77(12844)

? divided by 35 = 3 remainder24
x/35=3*35+24
x/35=105+24
x/35=129
x=129*35
x=4,515 ans.
Proof
4,515/35=3*35+24
129=105+24
129=129